OpenStudy (anonymous):

am i right? A store displays six computers on a shelf side–by–side. If the first computer is eight inches wide and each successive computer is four inches wider than the previous one, find the total width of the computers on the shelf. Part 1: Describe the sigma notation used in answering the question above. (2 points) Part 2: Show all your work and answer the question. (3 points)

OpenStudy (anonymous):

8 ∑ 4 + 4n n=1 ∑ 4 + 4n = 4 ∑ 1 + n 4 * 8 + 4 ∑ n 32 + 4 [8 * 9 / 2] 32 + 40 72 inches

jimthompson5910 (jim_thompson5910):

no, the answer is not \[\Large \sum_{n=1}^{8}4+4n\] for part 1 Part 2 depends on part 1, so that is also incorrect

OpenStudy (anonymous):

ok can you help me then @jim_thompson5910

jimthompson5910 (jim_thompson5910):

How many computers are we talking about here?

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

so you're not summing to 8 up at the top (of the sigma), you're summing to 6

OpenStudy (anonymous):

thats the problem above says so i guess so

jimthompson5910 (jim_thompson5910):

so it should be \[\Large \sum_{n = 1}^{6}4+4n\] for part 1

OpenStudy (anonymous):

so is the first and 2nd part for part 2 right?

jimthompson5910 (jim_thompson5910):

first part is, but the second is not...and you'd have to use 6 instead of 8

OpenStudy (anonymous):

∑ 4 + 4n = 4 ∑ 1 + n 4 * 6 + 4 ∑ n 24 + 4 [6 * 7 / 2] 24 + 25 49 inches

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

4 [6 * 7 / 2] doesn't become 25

OpenStudy (anonymous):

84? i added thats why

OpenStudy (anonymous):

108 inches ?

jimthompson5910 (jim_thompson5910):

it is 108 inches

jimthompson5910 (jim_thompson5910):

108 is the total final answer for part 2

OpenStudy (anonymous):

can you help me with another problem?

jimthompson5910 (jim_thompson5910):

sure, what is it

OpenStudy (anonymous):

help me with the steps i have an answer just don't know how? geometric partial sum ∑[i=1,5,3(-4)^(i-1)].

OpenStudy (anonymous):

nevermind i found it

jimthompson5910 (jim_thompson5910):

You're basically summing 3(-4)^(i-1) from i = 1 to i = 5 So add up 3(-4)^(i-1) five times like so 3(-4)^(i-1) 3(-4)^(i-1) 3(-4)^(i-1) 3(-4)^(i-1) 3(-4)^(i-1) Then replace the first i with 1, the second i with 2, the third with 3, etc etc like this 3(-4)^(1-1) 3(-4)^(2-1) 3(-4)^(3-1) 3(-4)^(4-1) 3(-4)^(5-1) Now evaluate

OpenStudy (anonymous):

but i do need help with What is the 5th partial sum of ∑[i=1,∞,-1+5n]

OpenStudy (anonymous):

the answer i got for the last problem was 11,170

jimthompson5910 (jim_thompson5910):

same thing, but you're now doing it with -1+5n

OpenStudy (anonymous):

i got 60 ?

jimthompson5910 (jim_thompson5910):

write out -1+5n five times -1+5n -1+5n -1+5n -1+5n -1+5n then replace the first n with 1, the second with 2, etc etc to get -1+5(1) -1+5(2) -1+5(3) -1+5(4) -1+5(5)

jimthompson5910 (jim_thompson5910):

Evaluate each piece, then add them all up

OpenStudy (anonymous):

70?

jimthompson5910 (jim_thompson5910):

yes, you are correct the answer is 70